OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2?

OpenStudy (anonymous):

2x – y = 5 2x – y = –5 2x + y = 5 2x + y = –5

OpenStudy (ghazi):

y+3=2(x-1)..now you can do it

OpenStudy (anonymous):

first you need the point slope formula \[y-y_1=m(x-x_1)\] with \[m=2,x_1=1,y_1=-3\]

OpenStudy (anonymous):

y-(-3)=2(x-1)

OpenStudy (anonymous):

y+3=2x-2

OpenStudy (anonymous):

2x+y=5..?